The length of rectangular garden exceeds three times the width by 12 meters. The area of the garden is 180m^2. Find the dimension of the garden
Answer by jorel555(1290)

Length=3xWidth+12. Let w equal width. Then:
w(3w+12)=180
3w^2+12w-180=0
w^2+4w-60=0
(w+10)(w-6)=0
w=6,-10
We can throw away the negative result, and we get a width of 6 meters, and a length of 30 meters!!!!!!!!!!!!!!!!!!!